OpenStudy (shaleiah):

A population of 170 crickets doubles in size every month. How many crickets will there be after 2 years?

OpenStudy (shaleiah):

@Michele_Laino

OpenStudy (mathmale):

Lot of crickets! Their count will have doubled 24 times (as there are 24 months in 2 years). What would be an appropriate formula for calculating the population of crickets after 24 months?

OpenStudy (mathmale):

To answer this, create an EXPONENTIAL

OpenStudy (shaleiah):

Alright. Mind giving me a moment?

OpenStudy (mathmale):

EQUATION and let n represent the number of times the population doubles. Have you done this kind of problem before? If so, type out the exponential growth model.

OpenStudy (shaleiah):

170*24^2=n

OpenStudy (shaleiah):

@Directrix

OpenStudy (mathmale):

The general formula for expo growth that I was looking for is\[y=a*e ^{kt}\]

OpenStudy (mathmale):

and we have to find the values of a and k from the given info. At the very beginning, t=0 and the population is 170. What is the value of the coefficient a?

Directrix (directrix):

Do you know the doubling formula? @shaleiah

Directrix (directrix):

You have it slightly off in your post: 170*24^2=n n = 170 ( (2) ^ 24 )

OpenStudy (shaleiah):

2,852,126,720

OpenStudy (shaleiah):

:D

OpenStudy (michele_laino):

hint: at first month the population is 170 at second month the population is 2*170 at the third month the population is \(2 \cdot 2 \cdot 170= 2^{(3-1)} \cdot 170\) at fourth month the population is \(2^{(4-1)}\cdot 170\) ... at the 24th month the population is \(2^{(24-1)}\cdot 170\)

Directrix (directrix):

I think the person starts with 170 crickets and has 340 at the end of month 1.

OpenStudy (mathmale):

I agree. However, it might be easier to find the value of the growth constant, k, in \[y=a*e ^{kt}\]

OpenStudy (mathmale):

If t=0, then y=a(1). But the initial value of y (initial population) is 170; therefore, a = 170. All you hae to do now is the find the value k of the growth constant.

OpenStudy (michele_laino):

please what are the corresponding options

OpenStudy (shaleiah):

Sorry, no options.

Directrix (directrix):

These may be they: a. 2,852,126,720 crickets c. 8,160 crickets b. 680 crickets d. 680 crickets

OpenStudy (michele_laino):

In my reasoning, I started to count from \(1\)

OpenStudy (michele_laino):

please I ask this: after 2 month what is the population?

OpenStudy (michele_laino):

months*

OpenStudy (michele_laino):

I think it is 170 at the first month and 170*2=340 at the second month

OpenStudy (michele_laino):

I think that the common sense suggests that reasoning